The total amount being spent on personal calls in a month by employees of a company follows a normal distribution with a mean of $900 and a standard deviation of $50 . Find the probability that in a randomly selected month the amount spent on personal calls is less than $750
i cant seem to figure this out or even know where to begin!... i would be so,so,so greatful if someone would point me in the right direction and at least give me some tools!please!

Answer by stanbon(75887)

The total amount being spent on personal calls in a month by employees of a company follows a normal distribution with a mean of $900 and a standard deviation of $50 . Find the probability that in a randomly selected month the amount spent on personal calls is less than $750
--------------------------------------------
Draw a normal curve; put u=900 in the middle and sigma = 50 under that.
Mark the 750 point on the horizontal axis(x-axis).
Shade the area under the curve to the left of 750.
That is the area you are looking for.
---
Draw a normal curve and let the horizontal axis be "z-scores".
Put 0 in the middle.
Find the z-score of 750:
z(750) = (750-900)/50 = -3
Mark -3 on the z-axis and shade the area to the left of z = -3.
-----------------------------
Now use a calculator, or software, or a z-chart to find the following:
P(x<750) = P(z<-3) = normalcdf(-100,-3) = 0.001350
-----------------------------
I used a TI calculator to find the probability.
Let me know if this "explanation" does not help.
-----------------------------
Cheers,
Stan .